# Learning from Reward-Free Offline Data: A Case for Planning with Latent Dynamics Models

### v3

@@ B.2.2 Evaluation
We evaluate agents in two scenarios: 1) How agents perform on test layouts when trained on various numbers of train layouts; 2) Given a constant number of training layouts, how agents perform on test maps with varying degrees of distribution shift from the training layouts.
@@
B.2.3 Results for Single Maze Setting
@@
Method Sucess rate)
@@
B.3 Ant U-Maze
To investigate whether our findings generalize to environments with more complicated control dynamics, we test the methods on the Ant U-Maze environment [TET12] with 8-dimensional action space and 29-dimensional state space. Similar to our previous analysis on Two-Rooms, we showcase the trajectory stitching capabilities of different methods.
B.3.1 Dataset Generation
To collect the dataset, we use a pretrained expert directional policy from [Par+24] to generate 55M transitions of exploratory data, where we resample a new direction every $10$ steps, and apply Gaussian noise with a standard deviation of $1.0$ to every action.
B.3.2 Evaluation
@@
Appendix C Models
For CRL, GCBC, GCIQL, and HIQL we use the implementations from the repository [^1] https://github.com/seohongpark/ogbench of OGBench [Par+24]. Likewise, for HILP we use the official implementation [^2] https://github.com/seohongpark/HILP from its authors.
For the Diverse PointMaze environment, to keep things consistent with our implementation of PLDM (C.1.4), instead of using frame stacking, we append the agent velocity directly to the encoder output.
C.1 PLDM
C.1.1 Objective for collapse prevention
To prevent collapse, we introduce a VICReg-based [BPL21] objective. We modify it to apply variance objective across the time dimension to encourage features to capture information that changes, as opposed to information that stays fixed [Sob+22]. The objective to prevent collapse is defined as follows:
$\mathcal{L}_{\mathrm{var}}=\frac{1}{HD}\sum^{H}_{t=0}\sum^{D}_{j=0}\mathrm{max}(0,\gamma-\sqrt{\mathrm{Var}(Z_{t,:,j})+\varepsilon})$
@@ C.1.1 Objective for collapse prevention
C.1.2 Ablations of Objective Components
@@
C.1.3 Model Details for Two-Rooms
We use the same Impala Small Encoder used by the other methods from OGBench [Par+24]. For predictor, we use the a 2-layer Gated Recurrent Unit [Cho14] with 512 hidden dimensions; the predictor input at timestep $t$ is a 2D displacement vector representing agent action at timestep $t$; while the initial hidden state is $h_{\theta}(s_{0})$, or the encoded state at timestep 0. A single layer normalization layer is applied to the encoder and predictor outputs across all timesteps. Parameter counts are the following:
1total params: 2218672
2encoder params: 1426096
3predictor params: 793600
C.1.4 Model Details for Diverse PointMaze Environment
@@
For the predictor input, we concatenate the state $s_{t}\in\mathbb{R}^{18\times 26\times 26}$ with the expanded action along the channel dimension. We denote action expansion with $\mathrm{Expander2D}(a_{t})\in\mathbb{R}^{2\times 26\times 26}$. The predictor output has the same dimension as the representation: $\hat{z}\in\mathbb{R}^{18\times 26\times 26}$. Both the encodings and predictions are flattened for computing the VicReg and IDM objectives.
@@
1total params: 53666
2encoder params: 33296
3predictor params: 20370
4
5PLDM(
6 (backbone): MeNet6(
7 (layers): Sequential(
8 (0): Conv2d(3, 16, kernel_size=(5, 5), stride=(1, 1))
9 (1): GroupNorm(4, 16, eps=1e-05, affine=True)
10 (2): ReLU()
11 (3): Conv2d(16, 32, kernel_size=(5, 5), stride=(2, 2))
12 (4): GroupNorm(8, 32, eps=1e-05, affine=True)
13 (5): ReLU()
14 (6): Conv2d(32, 32, kernel_size=(3, 3), stride=(1, 1))
15 (7): GroupNorm(8, 32, eps=1e-05, affine=True)
16 (8): ReLU()
17 (9): Conv2d(32, 32, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
18 (10): GroupNorm(8, 32, eps=1e-05, affine=True)
19 (11): ReLU()
20 (12): Conv2d(32, 16, kernel_size=(1, 1), stride=(1, 1))
21 )
22 (propio_encoder): Expander2D()
23 )
24 (predictor): ConvPredictor(
25 (layers): Sequential(
26 (0): Conv2d(20, 32, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
27 (1): GroupNorm(4, 32, eps=1e-05, affine=True)
28 (2): ReLU()
29 (3): Conv2d(32, 32, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
30 (4): GroupNorm(4, 32, eps=1e-05, affine=True)
31 (5): ReLU()
32 (6): Conv2d(32, 18, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
33 )
34 (action_encoder): Expander2D()
35 )
36)
C.1.5 Model Details for Ant-U-Maze
@@
1total params: 1080615
2encoder params: 9120
3predictor params: 1072007
4
5PLDM(
6 (backbone): MLPEncoder(
7 (globa_xy_encoder): Sequential(
8 (0): Linear(in_features=2, out_features=32, bias=True)
9 (1): LayerNorm((32,), eps=1e-05, elementwise_affine=True)
10 (2): Mish(inplace=True)
11 (3): Linear(in_features=32, out_features=256, bias=True)
12 (4): LayerNorm((256,), eps=1e-05, elementwise_affine=True)
13 )
14 (proprio_encoder): Identity()
15 )
16 (predictor): MLPPredictor(
17 (layers): Sequential(
18 (0): Linear(in_features=291, out_features=256, bias=True)
19 (1): LayerNorm((256,), eps=1e-05, elementwise_affine=True)
20 (2): Mish(inplace=True)
21 (3): Linear(in_features=256, out_features=256, bias=True)
22 (4): LayerNorm((256,), eps=1e-05, elementwise_affine=True)
23 (5): Mish(inplace=True)
24 (6): Linear(in_features=256, out_features=283, bias=True)
25 (7): LayerNorm((283,), eps=1e-05, elementwise_affine=True)
26 )
27 )
28)
Appendix D Effects of Uncertainty Regularization via Ensembles
@@
Appendix E Analyzing planning time of PLDM
@@
Table 5: Time of evaluation on one episode in Two-Rooms environment. Time is averaged across 25 episodes. PLDM success rates are normalized against the setting that replans every step. PLDM can match the latencies of GCIQL and HIQL by replanning less frequently with negligible cost to performance.
@@
Appendix F Latent Prediction vs. Input Reconstruction Learning
In this section, we evaluate reconstruction-based objectives for training the encoder and dynamics model. Specifically, we compare two approaches: one based on Dreamer [Haf+18], and one that replaces our VICReg objective with pixel reconstruction.
@@
We adapt DreamerV3 [Haf+23] to our setting by:
@@
  • Omit policy learning, instead planning via latent representation distance, just like we did for PLDM. Since the representations DreamerV3 uses are discrete, we use KL divergence instead of L2 distance.
@@
Results are shown in Table˜6. We test all methods on the two-rooms environment described in Section˜4.1. We use good-quality data, with long trajectories and good transition coverage. We find that pixel observation reconstruction is not a good objective to learn representations, and using it results in poor planning performance.
Table 6: Comparing reconstruction-based latent-dynamics learning to other baselines. We test the methods on good-quality data in the two-rooms environment. We see that reconstruction-based methods perform significantly worse than PLDM and other methods.
@@
Appendix G Analyzing HILP’s Out-of-Distribution Generalization
To understand HILP’s poor generalization to out-of-distribution (OOD) maze layouts, we visualize the distance in HILP’s learned latent representation space. HILP learns a latent representation $\phi(s)$, such that $\|\phi(s)-\phi(s_{g})\|_{2}$ is equal to the lowest number of transitions needed to traverse from $s$ to $s_{g}$. We hypothesize that $\phi$ fails to generalize to out-of-distribution maze layouts, resulting in incorrect predicted distances and in the failure of the goal-conditioned policy. We visualize the distances on in-distribution and out-of-distribution layouts for an encoder $\phi$ trained on 5 different layouts in Fig.˜9(b). We see that HILP distances are meaningful only on in-distribution layouts, and are very noisy on out-of-distribution layouts. This failure of the latent-space distance to generalize to out-of-distribution layouts confirms our hypothesis, and highlights the strength of PLDM.
(a) HILP distances on in-distribution maze layouts
(b) HILP distances on out-of-distribution maze layouts
Appendix H Analysing Statistical Significance of Results
In order to analyze whether the results in this paper are statistically significant, we perform Welch’s t-test to compare the performance of PLDM to other methods. Because 5 seeds is not enough for statistical tests, we pool results across settings and show results in Table˜7. We also run additional seeds for certain selected settings to get a total of 10 seeds per method, and show results of statistical analysis in Table˜8. Overall, we see that the results are significant, except for certain settings comparing to HILP, which aligns with our findings.
@@ Appendix H Analysing Statistical Significance of Results
CRL (p=2.64e-03) (p=6.26e-07) (p=1.27e-02) (p=3.43e-10) (p=9.15e-04)
GCBC (p=5.87e-04) (p=1.76e-08) (p=9.30e-03) (p=3.71e-20) (p=1.61e-05)
GCIQL (p=8.54e-03) (p=8.52e-08) (p=3.76e-02) (p=3.59e-08) (p=1.03e-03)
HILP (p=2.32e-02) (p=6.54e-06) (p=1.28e-05)
HIQL (p=2.26e-02) (p=6.40e-05) (p=3.33e-02) (p=1.69e-05) (p=4.09e-03)
@@
CRL (p=3.72e-10) (p=1.26e-07) (p=5.05e-09)
GCBC (p=1.77e-11) (p=7.02e-09) (p=8.26e-15)
GCIQL (p=6.87e-09) (p=6.64e-06) (p=4.72e-11)
HILP (p=2.55e-19)
HIQL (p=1.34e-09) (p=1.92e-06) (p=8.70e-10)
Appendix I Discussion of Computational Costs
@@
Unless listed below, all hyperparameters remain consistent with default values from OGBench1 and HILP2 repositories.
@@
For all methods, we used the learning rate of 3e-4. The rest of the hyperparameters were kept default.
Table 9: HILP hyperparameters
Hyperparam Value
Expectile 0.7
Skill expectile 0.7
Table 10: HIQL hyperparameters
Hyperparam Value
High-level AWR alpha 3.0
Low-level AWR alpha 3.0
Expectile 0.7
Table 11: GCIQL hyperparameters
Hyperparam Value
Actor-loss DDPG-BC
BC coefficient 0.3
Expectile 0.7
@@
Table 12: Dataset specific hyperparameters of CRL, GCBC, GCIQL, HIQL, HILP for the Diverse PointMaze environment. For HILP, we set the same value for expectile and skill expectile.
Dataset CRL GCBC GCIQL HIQL HILP
LR LR LR Expectile LR Expectile LR Expectile
# map layouts = 5 0.0003 0.0003 0.0002 0.8 0.0001 0.7 0.0001 0.9
# map layouts = 10 0.0003 0.0001 0.0001 0.9 0.0001 0.7 0.0001 0.9
# map layouts = 20 0.0003 0.0001 0.0001 0.6 0.0003 0.7 0.0001 0.9
# map layouts = 40 0.0003 0.0001 0.0003 0.9 0.0001 0.9 0.0001 0.9
@@
Planner $C_{\text{unc}}$ coeff $\beta$ 0.0001
Planner $C_{\text{unc}}$ coeff $\gamma$ 0.9
For the dataset specific hyperparameters, we tune the following parameters from Section˜C.1.1:
@@
Planner $C_{\text{unc}}$ coeff $\beta$ 1
Planner $C_{\text{unc}}$ coeff $\gamma$ 0.9
@@
Offline RL. This field aims to learn behaviors purely from offline data without online interactions. As opposed to imitation learning [Zar+24], offline RL is capable of learning policies that are better than the policy collecting the data. However, a big challenge is preventing the policy from selecting actions that were not seen in the dataset. CQL [Kum+20] relies on model conservatism to prevent the learned policy from being overly optimistic about trajectories not observed in the data. IQL [KNL21] introduces an objective that avoids evaluating the Q-function on state-action pairs not seen in the data to prevent value overestimation. MOPO [Yu+20] is a model-based approach to learning from offline data, and uses model disagreement to constrain the policy. See [Lev+20] for a more in-depth survey.
Foundation models in RL. Recently, following the success of NLP, the RL community put a lot of effort into training large sequence models, which sparked dataset collection efforts like Open-X-Embodiment [Col+23] and DROID [Kha+24]. Large datasets have enabled training models such as RT-2 [Bro+23] and Octo [Oct+24]. See [Yan+23a] for a more extensive survey on the topic.
Training representations for RL. Another way to use large amounts of data to improve RL agents is using self-supervised learning (SSL). CURL [LSA20] introduce an SSL objective in addition to the standard RL objectives. Later works also explore using a separate pre-training stage [Sch+21, Zha+22, Nai+22]. [Zho+24] show that pre-trained visual representations from DINO [Car+21, Oqu+23] can be used to learn a word model for planning.

### v4

@@ C.2.2 Evaluation
We evaluate agents in two scenarios: 1) How agents perform on test layouts when trained on various number of train layouts; 2) Given a constant number of training layouts, how agents perform on test maps with varying degrees of distribution shift from the training layouts.
@@
C.2.3 Results for Single Maze Setting
@@
Method Success rate
@@
C.3 Ant U-Maze
To investigate whether our findings generalize to environments with more complicated control dynamics, we test the methods on the Ant U-Maze environment [66] with 8-dimensional action space and 29-dimensional state space. Similar to our previous analysis on Two-Rooms, we showcase the trajectory stitching capabilities of different methods.
C.3.1 Dataset Generation
To collect the dataset, we use a pretrained expert directional policy from Park et al. [51] to generate 55M transitions of exploratory data, where we resample a new direction every $10$ steps, and apply Gaussian noise with a standard deviation of $1.0$ to every action.
C.3.2 Evaluation
@@
Appendix D Models
For CRL, GCBC, GCIQL, and HIQL we use the implementations from the repository [^1] https://github.com/seohongpark/ogbench of OGBench [51]. Likewise, for HILP we use the official implementation [^2] https://github.com/seohongpark/HILP from its authors.
For the Diverse PointMaze environment, to keep things consistent with our implementation of PLDM (D.1.4), instead of using frame stacking, we append the agent velocity directly to the encoder output.
D.1 PLDM
D.1.1 Objective for collapse prevention
To prevent collapse, we introduce a VICReg-based [4] objective. We modify it to apply variance objective across the time dimension to encourage features to capture information that changes, as opposed to information that stays fixed [61]. The objective to prevent collapse is defined as follows:
$\mathcal{L}_{\mathrm{var}}=\frac{1}{HD}\sum^{H}_{t=0}\sum^{D}_{j=0}\mathrm{max}(0,\gamma-\sqrt{\mathrm{Var}(Z_{t,:,j})+\epsilon})$
@@ D.1.1 Objective for collapse prevention
D.1.2 Ablations of Objective Components
@@
D.1.3 Model Details for Two-Rooms
We use the same Impala Small Encoder used by the other methods from OGBench [51]. For predictor, we use the a 2-layer Gated recurrent unit [14] with 512 hidden dimensions; the predictor input at timestep $t$ is a 2D displacement vector representing agent action at timestep $t$; while the initial hidden state is $h_{\theta}(s_{0})$, or the encoded state at timestep 0. A single layer normalization layer is applied to the encoder and predictor outputs across all timesteps. Parameter counts are the following:
total params: 2218672
encoder params: 1426096
predictor params: 793600
D.1.4 Model Details for Diverse PointMaze Environment
@@
For the predictor input, we concatenate the state $s_{t}\in\mathbb{R}^{18\times 26\times 26}$ with the expanded action $\mathrm{Expander2D}(a_{t})\in\mathbb{R}^{2\times 26\times 26}$ along the channel dimension. The predictor output has the same dimension as the representation: $\hat{z}\in\mathbb{R}^{18\times 26\times 26}$. Both the encodings and predictions are flattened for computing the VicReg and IDM objectives.
@@
total params: 53666
encoder params: 33296
predictor params: 20370
PLDM(
(backbone): MeNet6(
(layers): Sequential(
(0): Conv2d(3, 16, kernel_size=(5, 5), stride=(1, 1))
(1): GroupNorm(4, 16, eps=1e-05, affine=True)
(2): ReLU()
(3): Conv2d(16, 32, kernel_size=(5, 5), stride=(2, 2))
(4): GroupNorm(8, 32, eps=1e-05, affine=True)
(5): ReLU()
(6): Conv2d(32, 32, kernel_size=(3, 3), stride=(1, 1))
(7): GroupNorm(8, 32, eps=1e-05, affine=True)
(8): ReLU()
(9): Conv2d(32, 32, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
(10): GroupNorm(8, 32, eps=1e-05, affine=True)
(11): ReLU()
(12): Conv2d(32, 16, kernel_size=(1, 1), stride=(1, 1))
(propio_encoder): Expander2D()
(predictor): ConvPredictor(
(layers): Sequential(
(0): Conv2d(20, 32, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
(1): GroupNorm(4, 32, eps=1e-05, affine=True)
(2): ReLU()
(3): Conv2d(32, 32, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
(4): GroupNorm(4, 32, eps=1e-05, affine=True)
(5): ReLU()
(6): Conv2d(32, 18, kernel_size=(3, 3), stride=(1, 1), padding=(1, 1))
(action_encoder): Expander2D()
D.1.5 Model Details for Ant-U-Maze
@@
total params: 1080615
encoder params: 9120
predictor params: 1072007
PLDM(
(backbone): MLPEncoder(
(globa_xy_encoder): Sequential(
(0): Linear(in_features=2, out_features=32, bias=True)
(1): LayerNorm((32,), eps=1e-05, elementwise_affine=True)
(2): Mish(inplace=True)
(3): Linear(in_features=32, out_features=256, bias=True)
(4): LayerNorm((256,), eps=1e-05, elementwise_affine=True)
(proprio_encoder): Identity()
(predictor): MLPPredictor(
(layers): Sequential(
(0): Linear(in_features=291, out_features=256, bias=True)
(1): LayerNorm((256,), eps=1e-05, elementwise_affine=True)
(2): Mish(inplace=True)
(3): Linear(in_features=256, out_features=256, bias=True)
(4): LayerNorm((256,), eps=1e-05, elementwise_affine=True)
(5): Mish(inplace=True)
(6): Linear(in_features=256, out_features=283, bias=True)
(7): LayerNorm((283,), eps=1e-05, elementwise_affine=True)
Appendix E Effects of Uncertainty Regularization via Ensembles
@@
Appendix F Analyzing planning time of PLDM
@@
Table 5: Time of evaluation on one episode in Two-Rooms environment. Time is averaged across 25 episodes. PLDM success rates are normalized against the setting that replans every step. PLDM can match the latencies of GCIQL and HIQL by replanning less frequently with neligible cost to performance.
@@
F.1 Evaluating PLDM Performance With Adjusted Inference Compute Budget
As we see in Table˜5, replanning every 4 steps brings PLDM close to other methods’ inference speed. In order to further compare how PLDM compares to other methods when inference time compute is fixed between methods, we run additional evaluations, and report them in Figure˜10 and Figure˜11.
Figure 10: Comparing PLDM’s performance under fixed inference time compute budget on two-rooms. We see that across two-rooms experiments, PLDM performs only slightly worse when replanning every 4 steps compared to replanning every step.
Figure 11: Evaluating PLDM’s performance under fixed inference time compute budget on diverse maze. Similarly to two-rooms, we only see a minor degradation in performnace
Appendix G Extended Baselines: Input Reconstruction Learning and TD-MPC2
In this section, we evaluate reconstruction-based objectives for training the encoder and dynamics model, as well as TD-MPC2 [28]. For reconstruction, we compare two approaches: one based on Dreamer [25], and one that replaces our VICReg objective with pixel reconstruction.
@@ Appendix G Extended Baselines: Input Reconstruction Learning and TD-MPC2
We adapt DreamerV3 [26] to our setting by:
@@
  • Omitting policy learning, instead planning via latent representation distance, just like we did for PLDM. Since the representations DreamerV3 uses are discrete, we use KL divergence instead of L2 distance. We note that this is not what DreamerV3 architecture was designed for, making this comparison flawed. However, we believe it serves to highlight that compared to DreamerV3, PLDM is designed for a very different setting in terms of available data.
@@
TD-MPC2
Like Dreamer models, TD-MPC2 [28] also relies on rewards to learn its representions. In order to adapt TD-MPC2 to our setting, we remove the reward prediction component, and only keep the objective enforcing consistency between the predictor and encoder outputs.
Results are shown in Table˜6. We test all methods on the two-rooms environment described in Section˜4.1. We use good-quality data, with long trajectories and good transition coverage. We find that pixel observation reconstruction is not a good objective to learn representations, and using it results in poor planning performance. TD-MPC2 collapses altogether, achieving 0% success rate. To prevent collpase, we add inverse dynamics modelling (IDM), which somewhat improves performnace, although it is still far behind other methods. We only tested one seed for TD-MPC2 + IDM.
Table 6: Comparing reconstruction-based latent-dynamics learning and TDMPC-2 to other baselines. We test the methods on good-quality data in the two-rooms environment. We see that reconstruction-based methods perform significantly worse than PLDM and other baselines. TD-MPC2 fails to learn altogether due to collapse, and performs poorly even with the added IDM objective.
@@
TD-MPC2 110.0   ± 10.0
TD-MPC2 + IDM 135.0   ± 10.0
Appendix H Analyzing Statistical Significance of Results
In order to analyze whether the results in this paper are statistically significant, we perform Welch’s t-test to compare the performance of PLDM to other methods. Because 5 seeds is not enough for statistical tests, we pool results across settings and show results in Table˜7. We also run additional seeds for certain selected settings to get a total of 10 seeds per method, and show results of statistical analysis in Table˜8. Overall, we see that the results are significant, except for certain settings comparing to HILP and GCIQL, which aligns with our findings.
@@ Appendix H Analyzing Statistical Significance of Results
CRL (p=1.68e-03) (p=8.46e-07) (p=1.13e-02) (p=3.43e-10) (p=4.02e-03)
GCBC (p=4.25e-04) (p=1.46e-08) (p=7.26e-03) (p=3.71e-20) (p=1.94e-05)
GCIQL (p=1.29e-12) (p=1.22e-03)
HILP (p=2.11e-02) (p=6.54e-06) (p=4.10e-05)
HIQL (p=2.21e-02) (p=9.26e-05) (p=2.39e-02) (p=2.84e-06) (p=9.01e-03)
@@
CRL (p=3.82e-11) (p=1.13e-07) (p=6.76e-11)
GCBC (p=1.54e-12) (p=5.75e-10) (p=1.88e-14)
GCIQL
HILP (p=2.49e-19)
HIQL (p=8.89e-05) (p=2.83e-07) (p=2.65e-11)
Appendix I Analyzing HILP’s Out-of-Distribution Generalization
To understand HILP’s poor generalization to out-of-distribution (OOD) maze layouts, we visualize the distance in HILP’s learned latent representation space. HILP learns a latent representation $\phi(s)$, such that $\|\phi(s)-\phi(s_{g})\|_{2}$ is equal to the lowest number of transitions needed to traverse from $s$ to $s_{g}$. We hypothesize that $\phi$ fails to generalize to out-of-distribution maze layouts, resulting in incorrect predicted distances and in the failure of the goal-conditioned policy. We visualize the distances on in-distribution and out-of-distribution layouts for an encoder $\phi$ trained on 5 different layouts in Figure˜12(b). We see that HILP distances are meaningful only on in-distribution layouts, and are very noisy on out-of-distribution layouts. This failure of the latent-space distance to generalize to out-of-distribution layouts confirms our hypothesis, and highlights the strength of PLDM.
(a) HILP distances on in-distribution maze layouts
(b) HILP distances on out-of-distribution maze layouts
@@ Appendix I Analyzing HILP’s Out-of-Distribution Generalization
Unless listed below, all hyperparameters remain consistent with default values from OGBench1 and HILP2 repositories
@@
To tune hyperparameters, we ran 1 grid search for good quality data, then another grid search for sequence length 17, then tested both on all settings. In grid searches, we searched over learning rate, expectiles (for GCIQL, HILP, and HIQL), and the probability of sampling random goal. For CRL and GCBC, we ended up using default OGBench parameters. The other parameters are described below:
HILP.
We used learning rate of 3e-4, expectile of 0.7, and skill expectile of 0.7 in all settings.
GCIQL.
For settings with dataset size of 634 as well as for good quality data, we used expectile 0.7, learning rate 3e-4, BC coefficient of 0.3, and 0 probability of sampling random goal. For all other settings, we used learning rate 3e-4, expectile of 0.7, BC coefficient of 0.003, and probability of sampling a random goal 0.3.
HIQL.
For sequence length 17 and 33, we used learning rate of 3e-5, expectile of 0.7, AWR temperature of 3.0 (both high and low levels), and probability of sampling a random goal of 0.6. For all other settings, we used learning rate of 3e-4, expectile of 0.7, AWR temperature of 3.0 (both high and low levels), and probability of sampling a random goal of 0.
@@
Table 9: Dataset-specific hyperparameters of CRL, GCBC, and HILP for the Diverse PointMaze environment. For HILP, we set the same value for expectile and skill expectile.
Dataset CRL GCBC HILP
LR LR LR Expectile
5 layouts 0.0003 0.0003 0.0001 0.9
10 layouts 0.0003 0.0001 0.0001 0.9
20 layouts 0.0003 0.0001 0.0001 0.9
40 layouts 0.0003 0.0001 0.0001 0.9
Table 10: Dataset-specific hyperparameters of GCIQL and HIQL for the Diverse PointMaze environment.
Dataset GCIQL HIQL
LR Expectile BC Coeff. Prob Rand Goal LR Expectile AWR Temp. Prob Rand Goal
5 layouts 0.0003 0.6 0.3 0.15 0.00003 0.6 1.0 0.6
10 layouts 0.0003 0.6 0.3 0.15 0.0001 0.7 3.0 0.0
20 layouts 0.0003 0.6 0.3 0.15 0.00003 0.6 1.0 0.6
40 layouts 0.0003 0.6 0.3 0.15 0.0001 0.7 3.0 0.0
J.1.3 Ant U-Maze
Table 11: Ant-Umaze: dataset-specific hyperparameters for CRL, GCBC, and HILP, parsed from run names.
Seq Len CRL GCBC HILP
LR BC Coeff. Prob Rand Goal LR LR Expectile (skill / actor)
25 3e-5 0.1 0.3 3e-3 3e-4 0.6 / 0.8
50 3e-5 0.1 0.3 3e-3 3e-4 0.6 / 0.8
100 3e-4 0.1 0.0 3e-3 3e-4 0.6 / 0.8
250 3e-4 0.1 0.0 3e-3 3e-4 0.6 / 0.8
500 3e-4 0.1 0.0 3e-3 3e-4 0.6 / 0.8
Table 12: Ant-Umaze: dataset-specific hyperparameters for GCIQL and HIQL parsed from run names.
Seq Len GCIQL HIQL
LR Expectile BC Coeff. Prob Rand Goal LR Expectile AWR Temp. Prob Rand Goal
25 3e-5 0.9 0.15 0.15 3e-4 0.6 3.0 0.6
50 3e-5 0.9 0.15 0.15 3e-4 0.6 3.0 0.6
100 3e-5 0.9 0.15 0.15 3e-4 0.6 3.0 0.6
250 3e-4 0.9 0.3 0.0 3e-4 0.6 3.0 0.6
500 3e-4 0.9 0.3 0.0 3e-4 0.6 3.0 0.6
@@
Planner $C_{\text{uncertainty}}$ coeff $\beta$ 0.0001
Planner $C_{\text{uncertainty}}$ coeff $\gamma$ 0.9
For the dataset specific hyperparameters, we tune the following parameters from Section˜D.1.1:
@@
Planner $C_{\text{uncertainty}}$ coeff $\beta$ 1
Planner $C_{\text{uncertainty}}$ coeff $\gamma$ 0.9
@@
Offline RL. This field aims to learn behaviors purely from offline data without online interactions. As opposed to imitation learning [80], offline RL is capable of learning policies that are better than the policy collecting the data. However, a big challenge is preventing the policy from selecting actions that were not seen in the dataset. CQL [36] relies on model conservatism to prevent the learned policy from being overly optimistic about trajectories not observed in the data. IQL [33] introduces an objective that avoids evaluating the Q-function on state-action pairs not seen in the data to prevent value overestimation. MOPO [78] is a model-based approach to learning from offline data, and uses model disagreement to constrain the policy. See [41] for a more in-depth survey.
Foundation models in RL. Recently, following the success of NLP, the RL community put a lot of effort into training large sequence models, which sparked dataset collection efforts like Open-X-Embodiment [15] and DROID [31]. Large datasets have enabled training models such as RT-2 [10] and Octo [48]. See [75] for a more extensive survey on the topic.
Training representations for RL. Another way to use large amounts of data to improve RL agents is using self-supervised learning (SSL). CURL [38] introduce an SSL objective in addition to the standard RL objectives. Later works also explore using a separate pre-training stage [58, 82, 47]. Zhou et al. [83] show that pre-trained visual representations from DINO [13, 50] can be used to learn a word model for planning.
Appendix K Discussion of Computational Costs
Appendix L Limitations
All our experiments were conducted in navigation environments, excluding robot manipulation or partially observable settings. However, we argue that the conceptual understanding of the effects of data quality on the investigated methods will hold, as even in the relatively simple setting, we see many recent methods break down in surprising ways. Another limitation lies in the fact that PLDM is about 4 times slower during inference, although in Appendix˜F we find that even when limited to the same budget of inference compute as model-free methods, PLDM retains most of its performnace.